Calculate the equilibrium level of income if C=R100 million + 0,8y and I =R125 million
1
Expert's answer
2019-08-15T09:21:17-0400

The equilibrium income level is determined by the formula


Y = C + I


Y = 100 + 0.8Y + 125


(1-0.8) Y = (100 + 125)


0.2Y = 225


Y = 225 / 0.2


Y = 1125